5.2 The service time of the first service of a BMW is found to be normally distributed, with a mean of 70 minutes and a standard deviation of 9 minutes. Determine the following:
5.2.1 The probability that the first service will take more than an hour and a half.
5.2.2 The probability that the first service will take between 50 and 60 minutes.
i need urgent help. thank you
1
Expert's answer
2019-05-13T09:40:38-0400

If "X\sim N(\mu, \sigma^2)" then


"Z={X-\mu \over \sigma}\sim N(0, 1)"

Given that "\mu=70 \ min, \sigma=9 \ min."

5.2.1 The probability that the first service will take more than an hour and a half. 


"(1+1/2)\ h=90\ min"

"Z={90-70\over 9}={20\over 9}\approx 2.2222"

"P(X>90)=P(Z>{20\over 9})=1-P(Z\leq{20\over 9})="

"=1-0.986869=0.013131"

The probability that the first service will take more than an hour and a half is "0.013131."


5.2.2 The probability that the first service will take between 50 and 60 minutes. 


"Z_1={60-70\over 9}=-{10\over 9}\approx -1.1111"

"Z_2={50-70\over 9}=-{20\over 9}\approx -2.2222"

"P(50<X<60)=P({-20\over 9}<Z<{-10\over 9})="

"=P(Z<-1.1111)-P(Z<-2.2222)\approx"

"\approx 0.133260-0.013134=0.120126"

The probability that the first service will take between 50 and 60 minutes is "0.120126."